A man kicks a stationary ball of mass 200g, giving it a speed of 100cms. What impulse is imparted to the ball
1
Expert's answer
2019-11-25T10:08:42-0500

The impulse is


"I=mv"

"m=200\ g=0.2\ kg"

"v=100\frac{cm}{s}=1\frac{m}{s}"

"I=(0.2)(1)=0.2\frac{kgm}{s}"